A long, cylindrical wire of radius R has a current density J(r) = Jo(1 – r2/R2) for distances where r < R and J(r) = 0 for r < R where r is the distance from the center of the wire’s axis.

  1. Find the magnetic field strength inside (r < R) and outside (r > R) the wire.
  2. Sketch the magnetic field strength as a function of distance r from r = 0 to r = 2R.
  3. Find the location where the magnetic field strength is a maximum AND the value for B at that location.

Please be as detailed as possible when explaining the steps.
Answer #1

The magnetic field

According to Ampere Law 82R (inside) & Bide = lo s soda B & de = uo fto (1-22) (217846) r B (2018) = 460 Jo(212) 61-72.) radCC 2.ra ) = Jo (2Г) ( 4-82 / о B = Mo Jo / R — P 1e2 / а — до 15 — — в — Л. То 2R000 0 0 0 0 0 0 for maximum a (8-83 ) = 0 1-382 -0 2 R2 r2=282 2 R2 Ir=1 ₃ R Bmax = Mo Jo J R ( 1- 26 ) To 1 2 3R ( 2 ) Bmax